OpenStudy (anonymous):

Jack bought two sweaters for $17.99 and a pair slacks for $25.50. Including a 5 1/2% tax, how much was his total purchase?

OpenStudy (anonymous):

17.99 + 25.50 = 43.49 Tax = 43.49 x .055 = .................. s0 43.99 + tax = ................................total 6/

OpenStudy (anonymous):

what does 5 1/2 equal ?

OpenStudy (anonymous):

so 7.33 would be the answer if i divide it by 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my medal

OpenStudy (anonymous):

thanks :)

OpenStudy (phi):

what does 5 1/2 equal ? do you know how to change 1/2 to a decimal? 1/2 is 0.5 \[ 5 \frac{1}{2} \]menas 5 + 1/2 or 5 + 0.5 or just 5.5

OpenStudy (phi):

when you see percent as in 5.5% you should change it to a fraction \[ \frac{5.5}{100} \] or to a decimal 0.055

OpenStudy (phi):

the last step. If you want 5.5% of some number (for example 43.49) you first change 5.5% to 5.5/100 or 0.055 and then you multiply 0.055 times 43.49